OpenStudy (anonymous):

How does changing the function from f(x) = 3 sin 2x to g(x) = 3 sin 2x + 5 affect the range of the function?

OpenStudy (anonymous):

these are the answers The function shifts up 3 units, so the range changes from −1 to 1 in f(x) to 2 to 4 in g(x). The function shifts up 3 units, so the range changes from −3 to 3 in f(x) to 0 to 6 in g(x). The function shifts up 5 units, so the range changes from −1 to 1 in f(x) to 4 to 6 in g(x). The function shifts up 5 units, so the range changes from −3 to 3 in f(x) to 2 to 8 in g(x).

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (anonymous):

it shiftss up 5 units but id the rate

OpenStudy (sidsiddhartha):

range means the max and min values between which the function exists now tell me what are the max and min values of sinx

OpenStudy (anonymous):

how can i do that?

OpenStudy (sidsiddhartha):

maximum value of sinx is 1 which is sin90 and minimum value of sinx is -1 which is sin270 okay with this

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

so i can write \[-1\le sinx \le 1\] ok?

OpenStudy (anonymous):

yes :)

OpenStudy (sidsiddhartha):

this is also true for sin2x so i can also write \[-1\le \sin2x \le 1\]

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

now multiply 3 with each terms then it will be \[-3\le 3\sin2x \le 3\] ok??

OpenStudy (sidsiddhartha):

just multiplying 3 it wont affect all right!!

OpenStudy (sidsiddhartha):

any problem?

OpenStudy (anonymous):

so its The function shifts up 5 units, so the range changes from −3 to 3 in f(x) to 2 to 8 in g(x) bc we multiplied by 3?

OpenStudy (sidsiddhartha):

yeah so range of 3sin2x is -3 to 3 \[-3 \le 3\sin2x \le3\] now for range of 3sin2x+5 we have to add 5 with 3sin2x so we will have then \[-3+5\le 3\sin2x+5 \le 3+5\] \[2\le 3\sin2x+5 \le8\] so last option is correct :)

OpenStudy (anonymous):

ok thanks :)